OpenStudy (anonymous):

Determine weather (8.-9) is a solution of 4x-5y-3

hero (hero):

Hmm, what?

OpenStudy (anonymous):

Do you mean 4x-5y=3 (or -3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Substituting x=8 and y=-9 for x and y in the equation, we get 4(8)-5(-9)=3 32+45=3 but of course it does not. That is, that point (8,-9) does not balance the equation, and is therefor not a solution

jimthompson5910 (jim_thompson5910):

4x-5y=3 4(8)-5(-9)=3 32+45=3 77=3 Since the equation is NOT true, this means that (8,-9) is NOT a solution.